OpenStudy (jiteshmeghwal9):

One number is 7 more than the other and its square is 77 more than the square of the smaller number. What are the numbers?

OpenStudy (anonymous):

express the numbers in terms of x; \[x.... and..... 7 +x\]

OpenStudy (jiteshmeghwal9):

plz, give me the whole ans. step by step:)

OpenStudy (anonymous):

then: as the square of the smaller number is 77 less than the square of the bigger number we have this equation: \[(x +7) ^2 = (x^2) + 77\]

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (anonymous):

next expand: (x + 7)^2

OpenStudy (anonymous):

to give:\[x^2 + 2(7x) + 49\]

OpenStudy (anonymous):

then solve the equation, can you do it from here?

OpenStudy (jiteshmeghwal9):

I have a little confusion here:(

OpenStudy (jiteshmeghwal9):

where u expanded @bbsmilieface

OpenStudy (anonymous):

Let a and b be the largest and smallest numbers respectively. Solve the following for a and b:\[\left\{a=b+7,a^2=b^2+77\right\} \]\[\{a=9,b=2\} \]

OpenStudy (jiteshmeghwal9):

mean of expand?

OpenStudy (jiteshmeghwal9):

I can solve the eq. but didn't understand expand

OpenStudy (anonymous):

expansion of (x + 7) (x + 7): where the arrows are timesing